OpenStudy (anonymous):

Hey peeps (: Need a little help with this question for Simultaneous Equations. 3s square + 2t square = 11 and 3s + 2t = 1. Need a working too please (: Thanks!

OpenStudy (angela210793):

3s^2+2t^2=11 3s+2t=1--->s=(1-2t)/3 replace it in the 1st eq

OpenStudy (anonymous):

Ooh, okay (: Then t would be...? Could I have a full working too? Thanks!

OpenStudy (anonymous):

3s square means 3s^2 or 9s^2?

OpenStudy (anonymous):

3s^2 (:

OpenStudy (anonymous):

What will the full working be?

OpenStudy (anonymous):

U got the Wolfram link and a method, just follow it....

OpenStudy (anonymous):

Doesn't show workings tho :\

OpenStudy (anonymous):

Need some more help. All I need is just a full working right now. Thanks! (:

OpenStudy (anonymous):

U don't need help, u just need someone to do it all for u, u mean.

OpenStudy (anonymous):

... Erm, no, I've already done the working. Just need to compare it with some elses', that's all.. (:

OpenStudy (angela210793):

I guess Jimmy's solving it

OpenStudy (anonymous):

i'll start u off with the next step 3s^2+2t^2=11 s=(1-2t)/3 now plug this value of s into the first equation: 3 [(1-2t)/3]^2 + 2t^2 = 11 3 ( 1 - 4t + 4t^2) / 9 + 2t^2 - 11 = 0 (1/3)(1-4t+4t^2) + 2t^2 - 1 = 0 multiply through by 3 1 - 4t + 4t^2 + 6t^2 - 33 = 0 10t^2 - 4t - 32 = 0 5t^2 - 2t - 16 = 0 solve this equation to get t then substitute for t in the second of your original equatios to find s.

OpenStudy (anonymous):

So.. s = -1 and t = 2?

OpenStudy (angela210793):

u have to solve 5t^2 - 2t - 16 = 0 this 1st and from the values u'll get replace and see who makes the eq true